# Probabilistic Multi-Layer Perceptrons for Wind Farm Condition Monitoring

### v1

@@
Filippo Fiocchi Department of Computer Science, University College London, London, United Kingdom Domna Ladopoulou Department of Statistical Science, University College London, London, United Kingdom Petros Dellaportas Department of Statistical Science, University College London, London, United Kingdom Department of Statistics, Athens University of Economics and Business, Athens, Greece
@@
We provide a condition monitoring system for wind farms, based on normal behaviour modelling using a probabilistic multi-layer perceptron with transfer learning via fine-tuning. The model predicts the output power of the wind turbine under normal behaviour based on features retrieved from supervisory control and data acquisition (SCADA) systems. Its advantages are that (i) it can be trained with SCADA data of at least a few years, (ii) it can incorporate all SCADA data of all wind turbines in a wind farm as features, (iii) it assumes that the output power follows a normal density with heteroscedastic variance and (iv) it can predict the output of one wind turbine by borrowing strength from the data of all other wind turbines in a farm. Probabilistic guidelines for condition monitoring are given via a CUSUM control chart. We illustrate the performance of our model in a real SCADA data example which provides evidence that it outperforms other probabilistic prediction models.
@@ 1 Introduction
Due to the environmental and sustainability benefits, many energy producers have shifted their attention towards wind energy; one of the cleanest and fastest-growing sources of renewable energy. However, the severe weather conditions and the often remote locations of wind farms lead to both high operational failure rates and high maintenance costs that account for $20-30\%$ of the total cost related to power generation (Santolamazza et al.,, 2021). There is therefore a need to reduce the high operational and maintenance costs of wind turbines via a cost-effective and high-precision condition monitoring system. Such a system is crucial for an early fault detection and thus minimization of the prolonged wind turbine downtime.
@@
We develop a condition monitoring system based on a model that has all the four necessary characteristics described above based on a probabilistic multi-layer perceptron (PMLP) with transfer learning via fine-tuning. We assume that the output power follows a normal distribution with input-dependent mean and input-dependent variance which form our predictive density for each $10$-minute interval. For the probabilistic condition monitoring we propose the use of CUSUM control charts. We illustrate its performance in a real data application by comparing it with two other probabilistic models: a Gaussian process and a Bayesian neural network. We found that our model predicts better with respect to root mean square error, mean absolute error and maximum calibration error. We demonstrate that by transfer learning we can use information from all turbines in a wind farm to improve the prediction of the output power of a single wind turbine. We also provide a real data example in which our proposed condition monitoring system expressed via a CUSUM control chart reveals an early warning in a particular wind turbine failure.
@@ 2.1 Non-probabilistic normal behaviour modelling
A series of articles (Zaher et al.,, 2009; Schlechtingen and Santos,, 2011; Zhang and Wang,, 2014; Kusiak and Verma,, 2012; Bangalore and Tjernberg,, 2015; Xiang et al.,, 2021), proposed deep neural networks to model operational characteristics such as the temperature of the gearbox bearings, cooling oil and winding temperature of the wind turbine. Autoenconders have been proposed by Zhao et al., (2018) and Yang and Zhang, (2021) whereas generative adversarial networks combined with autoenconders have been used by Jin et al., (2021). Other methods include (Guo et al.,, 2012), who predicted the generator’s temperature with a nonlinear state estimate technique, and (Schlechtingen et al.,, 2013), who proposed an adaptive neuro-fuzzy interference system to predict the output power.
@@ 2.2 Probabilistic normal behaviour modelling
Due to their intrinsic probabilistic nature, Gaussian processes are common candidates for probabilistic normal behaviour modelling, see for example Pandit et al., (2020); Rogers et al., (2020); Li et al., (2014). The large impediment to their widespread application, namely their computational complexity which increases cubically with the data sample size, is usually being dealt with variational inference with inducing points (Titsias,, 2009). However, different input features increase the hyper-parameter dimension since proper implementation requires different kernels in each dimension, which creates both convergence and kernel identification problems. Finally, although the issue of heteroscedasticity can be solved by assuming that the error process also follows a Gaussian process (Lázaro-Gredilla and Titsias,, 2011), it is extremely difficult to resolve all three issues above simultaneously. Finally, although multi-task Gaussian processes is a modelling perspective that simultaneously exploits the data of all wind turbines, the computational complexity increases quickly and treating missing values from one or more wind turbines is not possible. The works of Lázaro-Gredilla and Titsias, (2011); Le et al., (2005) have incorporated noise heteroscedasticity whereas the trade-off between heteroscedasticity and computational complexity has been considered by Rogers et al., (2020). Recently, Deng et al., (2024) addressed the necessity of a probabilistic setting for power curve estimation by incorporating prior distributions on the weight parameters of a long short term memory (LSTM) Bayesian neural network. They improve their model by using a temporal convolutional neural network for temporal feature learning and an embedding layer to map discrete features (e.g., integer years) to dense vectors. The predictions are obtained via Monte Carlo sampling. Finally, Zhang et al., (2020) considered the problem of predicting output power conditional on future wind speed and direction forecasts and adopted a modelling approach that closely resembles ours. They employed an improved deep mixture density network by transforming the output energy to the interval $(0,1)$ and obtained probabilistic power predictions by assuming that it follows a mixture of Beta distributions. As is similarly emphasized in our study, they highlighted the fact that informed decision-making is one of the key benefits of a probabilistic approach to wind power prediction.
@@ 3.1 Probabilistic multi-layer perceptron
To model a heteroscedastic output noise we assume that for each $10$-minute interval the power output of a wind turbine $y$ conditioned on the SCADA input features $\mathbf{x}\in\Re^{d_{0}}$ follows a normal distribution with unknown mean and variance that both depend on $\mathbf{x}$, so $y\sim\mathcal{N}(\mu(\mathbf{x}),\sigma^{2}(\mathbf{x}))$. We will approximate $\mu(\mathbf{x})$ and $\sigma(\mathbf{x})$ by a probabilistic multi-layer perceptron (PMLP) with branching prediction heads. We now describe the architecture of our proposed model.
@@
where $\mathbf{W}_{i}\in\Re^{d_{i}\times d_{i-1}}$, $\mathbf{x}\in\Re^{d_{i-1}}$ and $\mathbf{b}_{i}\in\Re^{d_{i}}$. Furthermore, we adopt the Rectified Linear Unit function $ReLU_{i}:\Re^{d_{i}}\rightarrow\Re^{d_{i}}$ defined as
@@
for all $1\leq i\leq k^{*}$. Thus, before the $k^{*}$-th layer there is a common deep neural network; after the $k^{*}$-th layer, the network branches into two paths for predicting $\mu(\mathbf{x})$ and $\sigma(\mathbf{x})$ with each path having its own sequence of hidden layers with potentially different widths and depths. Note that the total sum of hidden layers is $k_{total}=k_{\mu}+k_{\sigma}-k^{*}$. A visual representation that corresponds to the choice of the PMLP architecture given $k_{\mu}=k_{\sigma}=4$, $k^{*}=3$, $d_{0}=34$, $d^{\mu}_{1}=d^{\sigma}_{1}=100$, $d^{\mu}_{2}=d^{\sigma}_{2}=80$, $d^{\mu}_{3}=d^{\sigma}_{3}=40$, $d^{\mu}_{4}=d^{\sigma}_{4}=20$ that was used in the application of Section 4 is depicted in Fig. 2. We call this particular PMLP architecture $A1$.
Figure 2: PMLP architecture A1 used in the application of Section 4 with shared initial layers and independent branches for mean and standard deviation.
@@ 3.2 Transfer learning via fine-tuning
Our proposed transfer learning is achieved via the following fine-tuning. We first train the LPMLP model (4) with the loss function (5). Then, we use the parameters of the pre-trained model as initial values, and we predict the mean and standard deviation of the output power of the wind turbine we are interested in. Thus, the training uses the loss function (3) that refers to one only wind turbine and is achieved very fast. Fig. 3 depicts the architecture of one LPMLP used in the application of Section 4 based on six wind turbines with $k_{\mu}=k_{\sigma}=3$, $k^{*}=3$, $d_{0}=34$, $d_{1}^{\sigma}=d_{1}^{\mu}=300$, $d_{2}^{\sigma}=d_{2}^{\mu}=200$, $d_{3}^{\sigma}=d_{3}^{\mu}=100$. We call this particular architecture A2. To show the scalability difference, we note that for our model in Fig. 3 the input vectors correspond to all features of all wind turbines, and the model has $91,002$ parameters whereas the PMPL model of Fig. 2 has only $16,508$ parameters.
Figure 3: LPMLP architecture A2 used for both pre-training on wind farm data and fine-tuning on wind turbine-specific data used in the application of Section 4 with shared initial layers.
For the advantages of transfer learning in similar problems see, for example, Ma et al., (2024). Notable theoretical and empirical motivations for introducing transfer learning in the given probabilistic regression context include (i) regularisation: pre-training the LPML allows learning general features and patterns from a larger amount of data which can be especially beneficial in the case where the operating conditions of the wind farm are extremely diverse. Thus, it is easier for the model to deal with conditions that have not been observed in the particular wind turbine under inspection; (ii) data sparsity: pre-training on the whole data of a wind farm dataset can help alleviate data sparsity by providing additional data for learning more general but not yet observed representations. For example, see Liao et al., (2023) for a discussion of limited data availability for newly built wind farms; and (iii) improved convergence: pre-training the LMPL facilitates faster converge during fine-tuning on a single wind turbine.
@@ 3.3 Condition Monitoring
Assume that the model training phase has ended and we are now in the live condition monitoring phase. To emphasize that the data is observed sequentially in time, we will be now using a subscript $t$ rather than $i$ for the time intervals. Assume that we observe $T$ consecutive power outputs $y_{t},\;t=1,2,\ldots,T$, and we have predictions from our model $N(\mu(\mathbf{x}_{t}),\sigma^{2}(\mathbf{x}_{t}))$. Under the hypothesis of model adequacy, or no-fault, we define $v_{t}=(y_{t}-\mu(\mathbf{x}_{t}))/\sigma(\mathbf{x}_{t}$) and test whether the sample $v_{t},\;t=1,2,\ldots,T$ comes from a $N(0,1)$ distribution. For example, we could perform a test on a daily basis for 10-minute intervals where $T=144$.
@@
where $S_{\mathrm{H}}(0)=S_{\mathrm{L}}(0)=0$. The reference (or allowance or slack) value of $k$ is traditionally specified with past experience or from simulations based on supervised settings. A common choice is about halfway between the target mean zero and the out-of-control value that we are interested in detecting quickly. Therefore, if we are interested in detecting a deviance of one standard deviation a plausible choice is $k=1/2$; see Montgomery, (2009). Recent advances propose an adaptive adjustment of its value (Wu et al.,, 2009). Note that $S_{\mathrm{H}}(t)$ and $S_{\mathrm{L}}(t)$ are viewed as cumulative deviations from zero that are greater than $k$ that are not allowed to take negative values. If either of them exceeds a decision interval $I$, the process is considered to be out-of-control. A reasonable choice of $I$ is five times the process standard deviation (Montgomery,, 2009) which is one in our case, so we propose $I=5$. For $t=1,\ldots,T$ the test statistics are defined as $A_{t}^{\mathrm{H}}=$ $\max\left\{S_{\mathrm{H}}(i),i=1,\ldots,t\right\}$ and $A_{t}^{\mathrm{L}}=\max\left\{S_{\mathrm{L}}(i),i=1,\ldots,T\right\}$. The null hypothesis of normal operation in the wind turbine is rejected (with a one-sided test) if $A_{n}=\max\left(A_{t}:1\leqslant t\leqslant n\right)$ exceeds its upper critical value, where $A_{t}=\max\left(A_{t}^{\mathrm{H}},A_{t}^{\mathrm{L}}\right)$.
@@ 4.1 Data Acquisition and Filtering
The application utilises $10$-minute SCADA and events data from the six Senvion MM92 wind turbines at Kelmarsh wind farm in the UK (Plumley,, 2022). The dataset spans from $3^{\text{rd}}$ January $2016$ to $1^{\text{st}}$ July $2021$ and comprises over $1.7$ million data points containing $110$ variables including date-time, wind speed, bearing temperature, and power output. The data includes $10$-minute averages, standard deviations, minimum and maximum values of all measured variables. The recording of wind speed summary statistics started at $25^{\text{th}}$ September $2017$ so we considered only data entries from this date onwards. An operational status and events file was utilised for data filtering to ensure consistent modelling of normal behaviour. These files provide valuable insights into the operational conditions of the wind turbines, covering a range of scenarios from technical failures to operational or environmental standbys and warnings. To enable accurate behaviour modelling, it was essential to remove out-of-control condition data records and base our model training solely on filtered data. The data elimination process is illustrated in Fig. 4 and 5 where plots of output powers against wind speeds are depicted for each wind turbine before and after the filtering.
@@
All our analyses used as input features $34$ operational and environmental variables are listed in Table 1. The final filtered dataset across the six wind turbines comprises $966,936$ data points, with $189,381$ of them belonging to one single wind turbine used in the application Section 4.2.
Table 1: SCADA operational and environmental variables used as input features in the proposed models. Avg: Average; Stdev: Standard deviation; Min: Minimum; Max: Maximum; temp: temperature.

### v2

@@
Filippo Fiocchi Department of Computer Science, University College London, London, United Kingdom Domniki Ladopoulou Department of Statistical Science, University College London, London, United Kingdom Corresponding author; e-mail: domna.ladopoulou.22@ucl.ac.uk Petros Dellaportas Department of Statistical Science, University College London, London, United Kingdom Department of Statistics, Athens University of Economics and Business, Athens, Greece
@@
We provide a condition monitoring system for wind farms, based on normal behaviour modelling using a probabilistic multi-layer perceptron with transfer learning via fine-tuning. The model predicts the output power of the wind turbine under normal behaviour based on features retrieved from supervisory control and data acquisition (SCADA) systems. Its advantages are that (i) it can be trained with SCADA data of at least a few years, (ii) it can incorporate all SCADA data of all wind turbines in a wind farm as features, (iii) it assumes that the output power follows a normal density with heteroscedastic variance and (iv) it can predict the output of one wind turbine by borrowing strength from the data of all other wind turbines in a farm. Probabilistic guidelines for condition monitoring are given via a cumulative sum (CUSUM) control chart, which is specifically designed based on a real-data classification exercise and, hence, is adapted to the needs of a wind farm. We illustrate the performance of our model in a real SCADA data example which provides evidence that it outperforms other probabilistic prediction models.
Abbreviation Definition
CP Coverage Probability
CUSUM Cumulative Sum
LPMLP Large Probabilistic Multi-Layer Perceptron
LSTM Long Short-Term Memory
MAE Mean Absolute Error
MCE Maximum Calibration Error
NMAE Normalized Mean Absolute Error
NRMSE Normalized Root Mean Square Error
PMLP Probabilistic Multi-Layer Perceptron
RBF Radial Basis Function
ReLU Rectified Linear Unit
RMSE Root Mean Square Error
SCADA Supervisory Control and Data Acquisition
@@ 1 Introduction
Due to the environmental and sustainability benefits, many energy producers have shifted their attention towards wind energy; one of the cleanest and fastest-growing sources of renewable energy. However, the severe weather conditions and the often remote locations of wind farms lead to both high operational failure rates and high maintenance costs that account for $20-30\%$ of the total cost related to power generation (Santolamazza et al.,, 2021). There is therefore a need to reduce the high operational and maintenance costs of wind turbines via a cost-effective and high-precision condition monitoring system. Such a system is crucial for an early fault detection and thus minimisation of the prolonged wind turbine downtime.
@@
We develop a condition monitoring system based on a model that has all the four necessary characteristics described above based on a probabilistic multi-layer perceptron (PMLP) with transfer learning via fine-tuning. We assume that the output power follows a normal distribution with input-dependent mean and input-dependent variance which form our predictive density for each $10$-minute interval. For the probabilistic condition monitoring we propose the use of cumulative sum (CUSUM) control charts. We illustrate its performance in a real-data application by comparing it with two other probabilistic models: a sparse Gaussian process and a Bayesian neural network, as shown in Table 2. We found that our model predicts better with respect to root mean square error (RMSE), mean absolute error (MAE) and maximum calibration error (MCE). We demonstrate that by transfer learning we can use information from all turbines in a wind farm to improve the prediction of the output power of a single wind turbine. We also provide a real data example in which our proposed condition monitoring system expressed via a CUSUM control chart reveals an early warning in a particular wind turbine failure.
@@ 2.1 Non-probabilistic normal behaviour modelling
A series of articles (Zaher et al.,, 2009; Schlechtingen and Santos,, 2011; Zhang and Wang,, 2014; Kusiak and Verma,, 2012; Bangalore and Tjernberg,, 2015; Xiang et al.,, 2021), proposed deep neural networks to model operational characteristics such as the temperature of the gearbox bearings, cooling oil and winding temperature of the wind turbine. Autoencoders have been proposed by Zhao et al., (2018) and Yang and Zhang, (2021) whereas generative adversarial networks combined with autoencoders have been used by Jin et al., (2021). Other methods include (Guo et al.,, 2012), who predicted the generator’s temperature with a nonlinear state estimate technique, and (Schlechtingen et al.,, 2013), who proposed an adaptive neuro-fuzzy inference system to predict the output power.
@@ 2.2 Probabilistic normal behaviour modelling
Due to their intrinsic probabilistic nature, Gaussian processes are common candidates for probabilistic normal behaviour modelling, see for example Pandit et al., (2020); Rogers et al., (2020); Li et al., (2014). The large impediment to their widespread application, namely their computational complexity which increases cubically with the data sample size, is usually being dealt with variational inference with inducing points (Titsias,, 2009). However, different input features increase the hyper-parameter dimension since proper implementation requires different kernels in each dimension, which creates both convergence and kernel identification problems. Finally, although the issue of heteroscedasticity can be solved by assuming that the error process also follows a Gaussian process (Lázaro-Gredilla and Titsias,, 2011), it is extremely difficult to resolve all three issues above simultaneously. Finally, although multi-task Gaussian processes is a modelling perspective that simultaneously exploits the data of all wind turbines, the computational complexity increases quickly and treating missing values from one or more wind turbines is not possible. The works of Lázaro-Gredilla and Titsias, (2011); Le et al., (2005) have incorporated noise heteroscedasticity whereas the trade-off between heteroscedasticity and computational complexity has been considered by Rogers et al., (2020). Recently, Deng et al., (2024) addressed the necessity of a probabilistic setting for power curve estimation by incorporating prior distributions on the weight parameters of a long short-term memory (LSTM) Bayesian neural network. They improve their model by using a temporal convolutional neural network for temporal feature learning and an embedding layer to map discrete features (e.g., integer years) to dense vectors. The predictions are obtained via Monte Carlo sampling. Finally, Zhang et al., (2020) considered the problem of predicting output power conditional on future wind speed and direction forecasts and adopted a modelling approach that closely resembles ours. They employed an improved deep mixture density network by transforming the output energy to the interval $(0,1)$ and obtained probabilistic power predictions by assuming that it follows a mixture of Beta distributions. As is similarly emphasized in our study, they highlighted the fact that informed decision-making is one of the key benefits of a probabilistic approach to wind power prediction.
@@ 3.1 Probabilistic multi-layer perceptron
To model a heteroscedastic output noise we assume that for each $10$-minute interval the power output of a wind turbine $y$ conditioned on the SCADA input features $\mathbf{x}\in\Re^{d_{0}}$ follows a normal distribution with unknown mean and variance that both depend on $\mathbf{x}$, so $y\sim\mathcal{N}(\mu(\mathbf{x}),\sigma^{2}(\mathbf{x}))$. We will approximate $\mu(\mathbf{x})$ and $\sigma(\mathbf{x})$ by a PMLP with branching prediction heads. We now describe the architecture of our proposed model.
@@
where $\mathbf{W}_{i}\in\Re^{d_{i}\times d_{i-1}}$, $\mathbf{x}\in\Re^{d_{i-1}}$ and $\mathbf{b}_{i}\in\Re^{d_{i}}$. Furthermore, we adopt the Rectified Linear Unit (ReLU) function $ReLU_{i}:\Re^{d_{i}}\rightarrow\Re^{d_{i}}$ defined as
@@
for all $1\leq i\leq k^{*}$. Thus, before the $k^{*}$-th layer there is a common deep neural network; after the $k^{*}$-th layer, the network branches into two paths for predicting $\mu(\mathbf{x})$ and $\sigma(\mathbf{x})$ with each path having its own sequence of hidden layers with potentially different widths and depths. Note that the total sum of hidden layers is $k_{total}=k_{\mu}+k_{\sigma}-k^{*}$.
In the application presented in Section 4, we compare two network architectures, A1 and A2. Architecture A1 is defined by $k_{\mu}=k_{\sigma}=4$, $k^{*}=3$, $d_{0}=41$, $d^{\mu}_{1}=d^{\sigma}_{1}=100$, $d^{\mu}_{2}=d^{\sigma}_{2}=80$, $d^{\mu}_{3}=d^{\sigma}_{3}=40$, $d^{\mu}_{4}=d^{\sigma}_{4}=20$, and architecture A2 by $k_{\mu}=k_{\sigma}=3$, $k^{*}=3$, $d_{0}=41$, $d_{1}^{\sigma}=d_{1}^{\mu}=300$, $d_{2}^{\sigma}=d_{2}^{\mu}=200$, $d_{3}^{\sigma}=d_{3}^{\mu}=100$. A visual representation of these architectures is provided in Figs. 2 and 3, respectively. We incrementally built these architectures by inspecting loss functions on training and testing data.
Figure 2: Architecture A1 used in the application of Section 4 for both PMLP and LPMPL with shared initial layers and independent branches for mean and standard deviation.
Figure 3: Architecture A2 used in the application of Section 4 for both PMLP and LPMPL with shared initial layers.
@@ 3.2 Transfer learning via fine-tuning
Our proposed transfer learning is achieved via the following fine-tuning. We first train the LPMLP model (4) with the loss function (5). Then, we use the parameters of the pre-trained model as initial values, and we predict the mean and standard deviation of the output power of the wind turbine we are interested in. Thus, the training uses the loss function (3) that refers to one only wind turbine and is achieved very fast. We apply this training method to both the A1 and A2 architectures to investigate whether additional parameters, combined with a pre-training phase, provide greater benefits to the model compared to pre-training alone. Specifically, architectures A1 and A2 differ in their branching and number of parameters, which are $16,508$ and $91,002$, respectively.
For the advantages of transfer learning in similar problems see, for example, Ma et al., (2024). Notable theoretical and empirical motivations for introducing transfer learning in the given probabilistic regression context include (i) regularisation: pre-training the LPMLP allows learning general features and patterns from a larger amount of data which can be especially beneficial in the case where the operating conditions of the wind farm are extremely diverse. Thus, it is easier for the model to deal with conditions that have not been observed in the particular wind turbine under inspection; (ii) data sparsity: pre-training on the whole data of a wind farm dataset can help alleviate data sparsity by providing additional data for learning more general but not yet observed representations. For example, see Liao et al., (2023) for a discussion of limited data availability for newly built wind farms; and (iii) improved convergence: pre-training the LMPLP facilitates faster converge during fine-tuning on a single wind turbine.
@@ 3.3 Condition Monitoring
Assume that the model training phase has ended and we are now in the live condition monitoring phase. To emphasize that the data is observed sequentially in time, we will be now using a subscript $t$ rather than $i$ for the time intervals. Assume that we observe $T$ consecutive power outputs $y_{t},\;t=1,2,\ldots,T$, and we have predictions from our model $N(\mu(\mathbf{x}_{t}),\sigma^{2}(\mathbf{x}_{t}))$. Under the hypothesis of model adequacy, or no-fault, we define $v_{t}=(y_{t}-\mu(\mathbf{x}_{t}))/\sigma(\mathbf{x}_{t}$) and test whether the sample $v_{t},\;t=1,2,\ldots,T$ comes from a $N(0,1)$ distribution. For example, we could perform a test over seventy-two hours with 10-minute intervals, where $T=432$.
@@
where $S_{\mathrm{H}}(0)=S_{\mathrm{L}}(0)=0$. The reference (or allowance or slack) value of $k$ is traditionally specified with past experience or from simulations based on supervised settings. A common choice is about halfway between the target mean zero and the out-of-control value that we are interested in detecting quickly. Therefore, if we are interested in detecting a deviance of one standard deviation a plausible choice is $k=1/2$; see Montgomery, (2009). Recent advances propose an adaptive adjustment of its value (Wu et al.,, 2009). Note that $S_{\mathrm{H}}(t)$ and $S_{\mathrm{L}}(t)$ are viewed as cumulative deviations from zero that are greater than $k$ that are not allowed to take negative values. If either of them exceeds a decision interval $I$, the process is considered to be out-of-control. A reasonable choice of $I$ is five times the process standard deviation (Montgomery,, 2009) which is one in our case, suggesting $I=5$. In practice, to define the decision interval $I$ according to the needs of the wind farm, $I$ can be estimated from empirical data to achieve the desired combination of recall and precision. In Section 4.3, we present an illustration in which we evaluate the model’s performance in terms of precision and recall. These metrics are defined as as $\textit{precision}=\textit{TP}/(\textit{TP}+\textit{FP})$ and $\textit{recall}=\textit{TP}/(\textit{TP}+\textit{FN})$ where TP, FP, and FN denote the number of true positives, false positives, and false negatives, respectively. For $t=1,\ldots,T$ the test statistics are defined as $A_{t}^{\mathrm{H}}=$ $\max\left\{S_{\mathrm{H}}(i),i=1,\ldots,t\right\}$ and $A_{t}^{\mathrm{L}}=\max\left\{S_{\mathrm{L}}(i),i=1,\ldots,T\right\}$. The null hypothesis of normal operation in the wind turbine is rejected (with a one-sided test) if $A_{n}=\max\left(A_{t}:1\leqslant t\leqslant n\right)$ exceeds its upper critical value, where $A_{t}=\max\left(A_{t}^{\mathrm{H}},A_{t}^{\mathrm{L}}\right)$.
@@ 4.1 Data Acquisition and Filtering
The application utilises $10$-minute SCADA and events data from the six Senvion MM92 wind turbines at Kelmarsh wind farm in the UK (Plumley,, 2022). The dataset spans from $3^{\text{rd}}$ January $2016$ to $1^{\text{st}}$ July $2021$ and comprises over $1.7$ million data points containing $110$ variables including date-time, wind speed, bearing temperature, and power output. The data includes $10$-minute averages, standard deviations, minimum and maximum values of all measured variables. The recording of wind speed summary statistics started at $25^{\text{th}}$ September $2017$ so we considered only data entries from this date onwards.
An operational status and events file was utilised for data filtering to ensure consistent modelling of normal behaviour. These files provide valuable insights into the operational conditions of the wind turbines, covering a range of scenarios from technical failures to operational or environmental standbys and warnings. To enable accurate behaviour modelling, it was essential to remove out-of-control condition data records and base our model training solely on filtered data. For filtering, periods of standby, warnings, and operational stops were excluded. Furthermore, data from the week preceding each forced outage was removed to reduce the likelihood of including out-of-control events in the training set. The data elimination process is illustrated in Fig. 4 and 5 where plots of output powers against wind speeds are depicted for each wind turbine before and after the filtering.
@@
All our analyses used as input features $41$ operational and environmental variables are listed in Table 1. The final filtered dataset across the six wind turbines comprises $846,968$ data points, with $163,562$ of them belonging to one single wind turbine used in the application Section 4.2.
Table 1: SCADA operational and environmental variables used as input features in the proposed models. Avg: Average; Stdev: Standard deviation; Min: Minimum; Max: Maximum; temp: temperature; dir: direction.
@@
Sine avg. wind speed dir. Sine max. wind speed dir.
Cosine avg. wind speed dir. Cosine max. wind speed dir.
Sine min. wind speed dir. Stdev. wind speed dir.
Cosine min. wind speed dir.